# Thinking Outside the (Gray) Box: A Context-Based Score for Assessing Value and Originality in Neural Text Generation

### v2

@@ Discussion
There are a few limitations worth noting. First of all, our score represents only a quantifiable approximation of a particular theoretical perspective on creativity, grounded in the dimensions of value and originality. It reflects a specific view of the evaluation of creativity based on the generated outputs and does not account for potential alternative theories (for example, arising from different cultures (Lubart 1999)) and perspectives (Rhodes 1961). In addition, our experiments are currently limited to only three relatively short-form text generation tasks. While their generalizability is supported by the theoretical framework discussed above, the resulting performance was experimentally evaluated for a finite number of scenarios.
Conclusion
In this paper, we presented CoVO, a novel score that quantifies the value and originality of neurally-generated text. The definition of CoVO is based on analysis of mutual information between the model’s outputs and inputs, and vice versa. We also proposed an optimization problem where a generative model aims to maximize this score to generate more creative products, and detailed how to use it in language modeling. We conducted experiments on poetry generation, math problem solving, and tasks included in NoveltyBench, exploring trade-offs in accuracy vs diversity. Effectively balancing value and originality maximization remains an open question, but our score seems to correlate with domain-specific measures appropriately. In addition, fine-tuning to maximize it improves quality- and diversity-related metrics.
Our research agenda aims to extend our method to other models and tasks, to include inference-level strategies such as creativity-oriented sampling schemes, and to explore its use for evaluation (e.g., in a Best-of-N setting (Stiennon et al. 2020)) rather than solely for optimization. We also plan to investigate the definition of additional scores for capturing other potentially relevant aspects of the creative process. Despite being costly and inherently constrained (Davis 2024), assessing whether our creativity score aligns with human judgment is another key direction for future work.
@@ Appendix A Implementation Details
Instead, the prompt used for computing $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ is:
@@
at inference time. Instead, for computing $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ we used the following:
@@
where user and assistant are keywords used by the model to identify different roles in the chat. Instead, for computing $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{y})$ we used the following:
@@ Appendix B GutenVerse Dataset
To evaluate the accidental reproduction rate of generated poems, we introduce the GutenVerse dataset, which comprises over 84,000 public-domain, English-written poems extracted from Project Gutenberg. While generated poems can reproduce different content, e.g., songs or copyrighted material, we believe this can provide a useful indication of how likely a text is original or not.
@@
We also released a datasheet (Gebru et al. 2018) for the GutenVerse dataset that can be found, together with the code used to create it, at: https://anonymous.4open.science/r/GutenVerse-DD32/.

### v3

@@ 7 Discussion
There are a few limitations worth noting. Firstly, our score represents only a quantifiable approximation of a particular theoretical perspective on creativity, grounded in the dimensions of value and originality. For example, value has been considered from the perspective of effectiveness, while other dimensions have been proposed as well (e.g., interestingness (Boden, 1994) or monetary worth (Lepak et al., 2007)); and other classic definitions of creativity, such as that presented in Boden (2003), add a third requirement by splitting originality into novelty and surprise. Moreover, our score reflects a specific view of the evaluation of creativity based on the generated outputs and does not account for potential alternative theories (for example, arising from different cultures (Lubart, 1999)) and perspectives (Rhodes, 1961). Finally, our experiments are currently limited to only three relatively short-form text generation tasks. While their generalizability is supported by the theoretical framework discussed above, the resulting performance was experimentally evaluated for a finite number of scenarios.
8 Conclusion
In this paper, we presented CoVO, a novel score that quantifies the value and originality of neurally-generated text. The definition of CoVO is based on the analysis of mutual information between the model’s outputs and inputs, and vice versa. We also proposed an optimization problem where a generative model aims to maximize this score to generate more creative products, and detailed how to use it in language modeling. We conducted experiments on poetry generation, math problem solving, and tasks included in NoveltyBench, exploring trade-offs in accuracy vs diversity. Effectively balancing value and originality maximization remains an open question, but our score seems to relate to domain-specific measures appropriately. In addition, fine-tuning to maximize it improves quality- and diversity-related metrics.
Our research agenda aims to extend our method to other models and tasks, to include inference-level strategies such as creativity-oriented sampling schemes, and to explore its use for evaluation (e.g., in a Best-of-N setting (Stiennon et al., 2020)) rather than solely for optimization. We also plan to investigate the definition of additional scores for capturing other potentially relevant aspects of the creative process. Despite being costly and inherently constrained (Davis, 2024), assessing whether our creativity score aligns with human judgment is another key direction for future work.
@@ Appendix A Implementation Details
Parameter Value
Total batches 100
Batch size $B$ 4
Gradient accumulation steps 8
Max new tokens 256
Temperature 1.
Top-$k$ 0
Optimizer Adam
Learning rate 1e-5
Max gradient normalization 100.
Rank (LoRA) 16
$\alpha$ parameter (LoRA) 32
Dropout (LoRA) 0.05
Training iterations 1
Scale rewards True
$\beta$ (when used) 0.05
Number of generations $G$ 4
Table 4: Training parameters for poetry generation.
@@
Instead, the prompt used for computing $p_{\bm{\theta}}(\mathbf{x}|\mathbf{y})$ is:
@@
at inference time. Instead, for computing $p_{\bm{\theta}}(\mathbf{x}|\mathbf{y})$ we used the following:
@@
where user and assistant are keywords used by the model to identify different roles in the chat. Instead, for computing $p_{\bm{\theta}}(\mathbf{x}|\mathbf{y})$ we used the following:
@@ Appendix B GutenVerse Dataset
To evaluate the accidental reproduction rate of generated poems, we introduce the GutenVerse dataset, which comprises over 84,000 public-domain, English-written poems extracted from Project Gutenberg. While generated poems can reproduce different content, e.g., songs or copyrighted material, we believe this can provide a useful indication of how likely a text is to be original or not.
@@
We also released a datasheet (Gebru et al., 2018) for the GutenVerse dataset that can be found, together with the code used to create it, at: https://anonymous.4open.science/r/GutenVerse-DD32/.